OpenStudy (anonymous):

WILL MEDAL Divide. -12/15 divided by (-2/5)= Express your answer in lowest terms.

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Here's the rule for dividing fractions: \(\sf\Large \frac{a}{b} \div \frac{c}{d} = \frac{a}{b} \times \frac{d}{c} = \frac{ad}{bc}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you do it with me. step by step

TheSmartOne (thesmartone):

ok first what is -1 x -1 = ?

OpenStudy (anonymous):

-12/15 divided by -2/5 then you flip -2/5 then you get -12/15 times 5/-2 do the rest

OpenStudy (anonymous):

just multiply across now

OpenStudy (anonymous):

ok... and thats the answer

OpenStudy (anonymous):